go online and find the recent annaual statement

(i cant find it please help me answer these questions)

Find the most recent annual financial statements for Coca Cola and do the following calculations. Be sure to label your answers (%, $, days, etc.) as applicable.

Return on Sales ____________

EBITDA _____________

Gross Profit Margin _______________

Operating Margin % _____________

Total Debt/Total Equity _____________

Total Debt/EBITDA _____________

Return on Assets _____________

Capital Expenditures _____________

Days Receivable _____________

EBITDA/Interest ______________

Fixed Charge coverage ______________

Return on Equity _______________

Days Inventory _______________

Current Ratio _______________
Answer #1

The follwing calculations and data is based on 2017 annual report of Coca Cola

Return on Sales=net income/sales=1248/35410=3.52%

EBITDA=Operating income (EBIT)+depreciation and amortization=7501+1260=8761 million

Gross profit margin=gross profit/sales=22154/35410=62.56%

Operating margin %=operating income/sales=7501/35410=21.18%

the above is answer..

we do only first 4 parts as per Chegg rule